Please help me solve for x:
x^2+4=y^2+6x-5

Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

x^2+4=y^2+6x-5
----------
x^2 - 6x = y^2 - 9
x^2 - 6x + 9 = y^2
(x-3)^2 = y^2
x-3 = y --> x = y+3
----
x-3 = -y --> x = -y+3
==

Answer by rothauserc(4718)

x^2 +4 = y^2 +6x -5
:
add 5 to from both sides of =
:
x^2 +9 = y^2 +6x
:
subtract 6x from both sides of =
:
x^2 -6x +9 = y^2
:
(x-3)^2 = y^2
:
take square root of both sides of =
:
x -3 = y
:
********
x = y +3
********
: